OpenStudy (anonymous):

graph the equation by plotting -3y=-x-6

OpenStudy (tkhunny):

You will need to pick some values. x from -2, -1, 0, 1, 2 might be a good place to start.

OpenStudy (anonymous):

can anyone please help me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-2,-1,0 for x is that ok

OpenStudy (tkhunny):

Sure. Substitute for x, each value successively and see what y-values is created.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(-2,-8) (-1,-7) (0,-6)

OpenStudy (anonymous):

is that right @tkhunny

OpenStudy (tkhunny):

No. Please try again. -3y=-x-6 x = -2 -3y = -(-2) - 6 -3y = 2 - 6 -3y = -4 y = 4/3 That's not very close to -8.

OpenStudy (anonymous):

I cant plot 4/3 it has to be a whole number

OpenStudy (tkhunny):

Well, if it HAS TO BE a whole number, you must select values carefully. You may wish to pick y-values and calculate corresponding x-values. I'll do the first. -3y=-x-6 y = -1 -3(-1) = -x - 6 3 = -x - 6 x + 3 = -6 x = -9 (-9,-1) I suggest more to the negative direction.

OpenStudy (anonymous):

-3 and -4 for x

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

You must pick values for y or you will get more fractions. I did y = -1 Try y = 0 Note: Forget what I said about "more negative". I was thinking about something else.

OpenStudy (anonymous):

0,-6 @tkhunny

OpenStudy (tkhunny):

You can't do that. You must show your work or I cannot help you. Just throwing out some final result simply is not constructive. -3y=-x-6 y = 0 -3(0) = -x - 6 0 = -x - 6 x = -6 (-6,0) This is the whole process, from beginning to end. Remember, we are picking y-values because you insisted. Try y = 1

OpenStudy (anonymous):

-3(1)=-x-6 -3=-x-6 x+3=-6 x=-3

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@tkhunny can you please help me

OpenStudy (tkhunny):

-3=-x-6 x+3=-6 No good. How did -3 turn into +3? x - 3 = -6

OpenStudy (anonymous):

would it be (1,-9)

OpenStudy (tkhunny):

You tell me. How did you get that? Did you add 3? -6 + 3 = -3

OpenStudy (anonymous):

i did this -6-3+-9

OpenStudy (anonymous):

i mean = -9

OpenStudy (tkhunny):

You had this x - 3 = -6 Why were you doing anything like -6-3+-9? You should just be adding 3.

OpenStudy (anonymous):

(-3,1)

OpenStudy (tkhunny):

You simply MUST get your algebra up to speed. It is very difficult to try to solve a given problem without reliable algebra skills. Practice, practice...